Nine years more than twice Bob's age is 5 years more than three times Harry's age. Find their ages.
Found 2 solutions by lynnlo, Alan3354:
Answer by lynnlo(4176)
Answer by Alan3354(69443)

Nine years more than twice Bob's age is 5 years more than three times Harry's age. Find their ages.
--------------------
9 + 2*Bob = 5 + 3*Harry
2B + 4 = 3H
If you assume only positive integer solutions,
Bob is 1, Harry is 2
Bob is 4, Harry is 4
Bob is 7, Harry is 6
etc
Bob's age = 3n+1, n = 0,1,2,3...
Harry's age = (2B + 4)/3